OpenStudy (anonymous):

solve 5x^2+13x=6 a.-2/5, 3 b.-3, 5/2 c.-5/2, 3 d0-3, 2/5

OpenStudy (anonymous):

tahts not an answer

hero (hero):

5x^2+13x=6 5x^2+13x - 6 = 0 5x^2 +15x-2x - 6 = 0 5x(x+3)-2(x+3) = 0 (x+3)(5x-2) = 0 x = -3, 2/5